OpenStudy (anonymous):

The ratio of boys to girls in class 7M is exactly 5:7. A new girl joins the class. The ratio of boys to girls is now 2:3. Find the number of boys and girls in the class now?

OpenStudy (anonymous):

The number of girls i gonna ba an even number and the number of boys is in the 3 times tables

OpenStudy (anonymous):

Say, there are an x amount of boys. Then: 7x=5y When the girl joins, it's: 3x=2(y+1) 1.5x=y+1 1.5x-1=y 7x=5*(1.5x-1) 7x=7.5x-5 -0.5x=-5 x=10 y=7*10/5=14 14+1=15 There are 10 boys and 15 girls.